# Surface topological quantum criticality: Conformal manifolds and Discrete Strong Coupling Fixed Points

### v1

@@ VII.2 Conformal manifolds as coset space $G/H$
Let us write down the matrix form of $\hat{V}_{c}$ when $\lambda_{3}=\epsilon$, $\lambda_{1}=\lambda_{2}=0$ is,
@@
Since the two diagonal entries are zero, the invariant subgroup is $SO(2)$. Therefore the coset space is,
@@
The matrix form of $\hat{V}_{c}$ when $\lambda_{3}=0$, $\lambda_{1}=\lambda_{2}=\epsilon$ is,
@@
Table 4: Table showing the list of isolated fixed points/fixed point manifolds, the shape of the manifold in the parameter space, and their stability for the $\mathcal{N}=3$ flavor surface. The coefficients $V^{c}_{i}$’s ($i=0,1,2,3$) are defined in Eqn.73. $\hat{V}^{c}$ is the interaction matrix at the fixed point/manifold. $\theta$ and $\phi$ are the polar and azimuthal angles respectively.
@@
VII.3 General solution: Triad basis
@@
To understand why $\chi$ is redundant, let us consider some $SO(3)$ rotations of the interaction matrix $\hat{V}$ defined in Eqn.73 carefully. If we rotate $\hat{V}$ by varying $\theta$ and $\phi$, we are indeed taken to different points on the manifold. Since the 4-dimensional space we defined in Eqn.73 is exclusively for a given point on the 2-sphere, as anticipated, these particular rotations effectively take the $\hat{V}$ matrix out of the 4-dimensional space defined for a given point in the conformal manifold.
@@
Plugging this back to the RG equation in Eqn.27, we arrive at the following set of dynamical equations for the four variables,
@@
We examine the fixed points of the theory. Consistent with the analysis of the RG flow of eigenvalues in section VIIA, we identify four distinct classes of fixed points, each characterized by its trace. Note that the coefficient of $\hat{T}_{0}$ is proportional to the trace. Table. 4 lists all the fixed points/manifolds and their properties. Once we solve for $V^{c}_{i}$’s, we then plug them back into the Eqn.73 to get the general matrix structure of the fixed point manifolds. For the weak and strong coupling fixed points, we find the fixed point interaction matrices are,
@@
VII.4 Properties of the fixed points and manifolds
@@
Having found the general solutions for the fixed points/manifolds, let us study the stability properties of each of them. This can be studied by linearizing the beta function in Eqn.27 around each fixed point. The properties of the weak and the strong coupling fixed points are similar to that of the $\mathcal{N}=2$ theory discussed in section VI. The weak coupling point is stable (all six operators are irrelevant), while the strong coupling fixed point is highly unstable (all six operators are relevant). These two are the isolated fixed points of the theory. In the following, we shall examine the properties of the two 2-sphere conformal manifolds in the parameter space.
VII.4.1 Conformal manifold $\hat{V}^{c}=\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$
@@
Let us study the conformal manifold defined by $\hat{V}^{c}=\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$. This is the one with the trace equal to unity shown in table.3. The eigenvalues at any point on the 2-sphere are $(\lambda_{1},\lambda_{2},\lambda_{3})=(\epsilon,0,0)$ and independent of both $\theta$ and $\phi$. So the fixed point shown in the table.3 is a point on this 2-sphere manifold.
@@
Analyzing the beta function in Eqn.27 around any point on this manifold reveals three irrelevant operators and one relevant operator. That is
$n_{IRR}=3,\,n_{MAR}=2\,\text{and}\,\,n_{R}=1.$ (78)
@@
Therefore, there is a three-dimensional subspace composed entirely of irrelevant operators in the four-dimensional space orthogonal to the tangent space at any point on the sphere. Consequently, if the bare interaction matrix elements lie within this three-dimensional subspace, the theory will flow towards the conformal 2-sphere manifold in the infrared limit.
@@
However, there is one relevant operator in the six-dimensional parameter space. To achieve quantum criticality, we must fine-tune this relevant operator to the fixed point manifold. Given that only one of the six operators requires tuning, we can infer that this conformal manifold is infrared-stable. Appendix C lists the marginal, relevant, and irrelevant operators and their respective scaling dimensions for a few selected points on the conformal manifold for reference.
VII.4.2 Conformal manifold $\hat{V}^{c}=\epsilon\hat{I}-\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$
@@
The conformal manifold defined by $\hat{V}^{c}=\epsilon\hat{I}-\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$ correspond to the one with trace equal to two in the table.3. The eigenvalues remain the same throughout the manifold, and they are given by $(\lambda_{1},\lambda_{2},\lambda_{3})=(\epsilon,\epsilon,0)$. Repeating the same process as in the previous case, we find that at any point on the manifold, there are three relevant operators and only one irrelevant operator, all lying in the four-dimensional space normal to the tangent space of two marginal operators. That is
$n_{IRR}=1,\,n_{MAR}=2\,\text{and}\,\,n_{R}=3.$ (79)
@@
Overall three operators must be fine-tuned to the fixed point for the system to exhibit quantum critical behavior in the infrared limit. Hence, this conformal manifold is infrared-unstable. See appendix C for the exact form of the relevant and irrelevant operators at a few selected points on the conformal manifold.
VII.5 Phase boundary
@@
Recall that one of the conformal manifolds defined by $\hat{V}^{c}=\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$ is infrared stable. It has only one relevant operator, as shown in Eqn.78 (also in Table.4). Therefore, this conformal manifold is naturally embedded within the phase boundary of the gapless state. Eqn.78 suggests there are three irrelevant operators ($n_{IRR}=3$) at any point on the manifold. Following the spirit of the analysis in the $\mathcal{N}=2$ case, we can illustrate that locally the phase boundary shall be made up of this 3-dimensional subspace of irrelevant flows plus the 2-dimensional 2-sphere conformal manifold. To summarize, we shall have a five-dimensional space as the phase boundary. The sole relevant operator is directed orthogonal to this 5-dimensional phase boundary.
@@
VII.6 Effective Yukawa field theory for the 5-dimensional phase boundary
In the previous section, where we studied the quantum criticality of the $\mathcal{N}=2$ phase, we developed an effective single boson Yukawa field theory for the conformal manifold of the shape of a ring in the three-dimensional parameter space (see section VIC). We also found that the theory works well for the phase boundary of the gapless surface state. The two main ingredients that helped us develop this EFT are 1) the interaction matrix at the ring manifold is factorizable, and 2) there is only one relevant direction out of the conformal manifold that could be associated with the boson mass.
We observe that the phase boundary of the $\mathcal{N}=3$ flavor TI surface also satisfies the two ingredients. First, the 2-sphere conformal manifold embedded within the phase boundary is also factorizable. It is evident from its expression, $\hat{V}_{c}=\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$, already in a tensor product form. Second, Eqn.78 suggests that the manifold has only one relevant operator. Therefore, the phase boundary of the $\mathcal{N}=3$ flavor TI surface near the infrared stable conformal manifold can also be expressed by a single boson Yukawa field theory in the infrared limit.
@@ VII.6 Effective Yukawa field theory for the 5-dimensional phase boundary
Figure 9: RG flows of the parameters $V_{0}$, $V_{z}$ and $V_{x}$ when three-loop effects are taken into account for an $\mathcal{N}=2$ TI surface. The smooth conformal manifold is defined by the Eqn.85. The perfect ring manifold shown in Fig.7 is deformed but the closed loop structure remains intact. The definitions of other RG flow lines are the same as in fig.7.
VIII Effects of $N_{L}$-loops, $N_{L}=2,3,4,…$
Further inclusion of effects of multiple-loop diagrams in the renormalization group equations can lead to the breaking of the emergent $SO(\mathcal{N})$ symmetry in the case of $\mathcal{N}=2,3$ studied in Section VI,VII. The number of loops $N_{L}$ is related to the number of four fermion vertices $N_{v}$. $N_{v}=N_{L}+1$. In the context of the $\epsilon$-expansion, these contributions with multiple loops are also represented as higher-order effects so that symmetry is only weakly broken. For instance, the $N_{L}$-loop effects are proportional to $\epsilon^{L+1}$, $L\geq 1$. For two-loop effects, the contributions are of the order of $\epsilon^{3}$, while for three-loop effects, the contributions are of the order of $\epsilon^{4}$.
@@ VIII Effects of $N_{L}$-loops, $N_{L}=2,3,4,…$
where $\beta(\hat{V})$ represents all the one-loop effects that already have been studied while $\beta_{M}(\hat{V})$ includes other multiple-loop effects that were not included in the studies in section VI.
@@
VIII.1 A three-loop effect of Type A: Distorted Conformal Manifolds
@@
Let us see explicitly if the one-dimensional conformal manifold exists when the three-loop effects are included in the RG equation. Let us express the interaction matrix $\hat{V}$ in terms of Pauli matrices as shown in Eqn.31. The RG equation is then (See fig.6 for the Feynman diagram),
@@
Table 5: Table showing the list of fixed points and their stabilities for the $\mathcal{N}=2$ flavor surface when two-loop effects are included.
@@
VIII.2 A two-loop effect of type B: Inducing flow in the Manifold
@@
Upon including the field renormalization effects, the RGE expressed in terms of $V_{0},V_{x},V_{z}$(see Eqn.31) takes the form,
@@
where $a=\frac{1}{32}$.
VIII.2.1 The fate of the conformal manifolds
@@
Taking into account the last two equations in Eqn.88, one can show that a finite flow is induced around the ring given by,
@@
Table.5 lists the fixed points of the RGE in Eqn.88 . Fig.10 shows the RG flow lines of the solutions to Eqn.88. Appendix.D lists the relevant and irrelevant operators and their scaling dimensions.

### v3

@@ VIII.2 Conformal manifolds as coset space $G/H$
Let us write down the matrix form of $\hat{V}_{c}$ when $(\lambda_{1},\lambda_{2},\lambda_{3})=(0,0,\epsilon)$ is,
@@
Since the two diagonal entries are zero, the invariant subgroup is $SO(2)$. Therefore, the coset space is,
@@
The matrix form of $\hat{V}_{c}$ when $(\lambda_{1},\lambda_{2},\lambda_{3})=(\epsilon,\epsilon,0)$ is,
@@
Table 4: Table showing the list of isolated fixed points/fixed point manifolds, the shape of the manifold in the parameter space, and their stability for the $\mathcal{N}=3$ flavor surface. The coefficients $V^{c}_{i}$’s ($i=0,1,2,3$) are defined in Eqn.76. $\hat{V}^{c}$ is the interaction matrix at the fixed point/manifold. $\theta$ and $\phi$ are the polar and azimuthal angles respectively. Also, $\epsilon=D-2$, where $D$ stands for spacetime dimensions.
@@
VIII.3 General solution: Triad basis
@@
To understand why $\chi$ is redundant, let us consider some $SO(3)$ rotations of the interaction matrix $\hat{V}$ defined in Eqn.76 carefully. If we rotate $\hat{V}$ by varying $\theta$ and $\phi$, we are indeed taken to different points on the manifold. Since the 4-dimensional space we defined in Eqn.76 is exclusively for a given point on the 2-sphere, as anticipated, these particular rotations effectively take the $\hat{V}$ matrix out of the 4-dimensional space defined for a given point in the conformal manifold.
@@
Plugging this back to the RG equation in Eqn.28, we arrive at the following set of dynamical equations for the four variables,
@@
We examine the fixed points of the theory. Consistent with the analysis of the RG flow of eigenvalues in section VIIIA, we identify four distinct classes of fixed points, each characterized by its trace. Note that the coefficient of $\hat{T}_{0}$ is proportional to the trace. Table. 4 lists all the fixed points/manifolds and their properties. Once we solve for $V^{c}_{i}$’s, we then plug them back into the Eqn.76 to get the general matrix structure of the fixed point manifolds. For the weak and strong coupling fixed points, we find the fixed point interaction matrices are,
@@
VIII.4 Properties of the fixed points and manifolds
@@
Having found the general solutions for the fixed points/manifolds, let us study the stability properties of each of them. This can be studied by linearizing the beta function in Eqn.28 around each fixed point. The properties of the weak and the strong coupling fixed points are similar to that of the $\mathcal{N}=2$ theory discussed in section VII. The weak coupling point is stable (all six operators are irrelevant), while the strong coupling fixed point is highly unstable (all six operators are relevant). These two are the isolated fixed points of the theory. In the following, we shall examine the properties of the two 2-sphere conformal manifolds in the parameter space.
VIII.4.1 Conformal manifold $\hat{V}^{c}=\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$
@@
Let us study the conformal manifold defined by $\hat{V}^{c}=\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$. This is the one with the trace equal to unity shown in table.3. The eigenvalues at any point on the 2-sphere are $(\lambda_{1},\lambda_{2},\lambda_{3})=(0,0,\epsilon)$ and independent of both $\theta$ and $\phi$. So the fixed point shown in the table.3 is a point on this 2-sphere manifold.
There exist two marginal operators that span the 2-dimensional tangent space on the conformal manifold, i.e., $n_{MAR}=2$. Let $\mathcal{O}^{\theta,\phi}_{M,1}(x)$, $\mathcal{O}^{\theta,\phi}_{M,2}(x)$ denote the two marginal operators. At a point $(\theta,\phi)$ on the manifold, we have,
$\mathcal{O}^{\theta,\phi}_{M,1}(x)$ $=$ $\sin 2\theta\cos^{2}\phi\,\Phi^{\dagger}_{1}\Phi_{1}+\frac{\sin 2\theta\sin 2\phi}{2}\Phi^{\dagger}_{1}\Phi_{2}+\cos 2\theta\cos\phi\Phi^{\dagger}_{1}\Phi_{3}+\sin 2\theta\sin^{2}\phi\,\Phi^{\dagger}_{2}\Phi_{2}$
[…]
$\mathcal{O}^{\theta,\phi}_{M,2}(x)$ $=$ $-\sin\theta\sin 2\phi\Phi^{\dagger}_{1}\Phi_{1}+\sin\theta\cos 2\phi\,\Phi^{\dagger}_{1}\Phi_{2}-\cos\theta\sin\phi\Phi^{\dagger}_{1}\Phi_{3}+\sin\theta\sin 2\phi\,\Phi^{\dagger}_{2}\Phi_{2}$ (81)
[…]
where $\Phi_{i}(x)=\psi^{T}_{i}is_{y}\psi_{i}(x)$ is the Cooper pair annihilation operator of flavor $i$($i=1,2,3$). As discussed in the section.III, the scaling dimensions of the two marginal operators, $\triangle_{\mathcal{O}_{M,1(2)}}=D$, the space-time dimensions.
@@
Analyzing the beta function in Eqn.28 around any point on this manifold reveals three irrelevant operators and one relevant operator. That is
$n_{IRR}=3,\,\text{and}\,\,n_{R}=1.$ (82)
@@
Therefore, there is a three-dimensional subspace composed entirely of irrelevant operators in the four-dimensional space orthogonal to the tangent space of the manifold. Consequently, if the bare interaction matrix elements lie within this three-dimensional subspace, the theory will flow towards the conformal 2-sphere manifold in the infrared limit.
@@
However, there is one relevant operator in the six-dimensional parameter space. To achieve quantum criticality, we must fine-tune this relevant operator to the fixed point manifold. Given that only one of the six operators requires tuning, we can infer that this conformal manifold is infrared-stable. Appendix D lists the marginal, relevant, and irrelevant operators and their respective scaling dimensions for a few selected points on the conformal manifold for reference.
VIII.4.2 Conformal manifold $\hat{V}^{c}=\epsilon\hat{I}-\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$
@@
The conformal manifold defined by $\hat{V}^{c}=\epsilon\hat{I}-\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$ correspond to the one with trace equal to $2\epsilon$ in the table.3. The eigenvalues remain the same throughout the manifold, and they are given by $(\lambda_{1},\lambda_{2},\lambda_{3})=(\epsilon,\epsilon,0)$.
Interestingly, we notice here that the functional form of the two marginal operators of this conformal manifold is the same as $\mathcal{O}^{\theta,\phi}_{M,1}(x)$ and $\mathcal{O}^{\theta,\phi}_{M,2}(x)$ ( Eqn.81), the marginal operators of the previous one. However, note that the two manifolds occupy distinct locations within the parameter space. Thus, while the marginal operators appear structurally identical in both cases, they are defined at different points in the parameter space.
Repeating the same process as in the previous case, we find that at any point on the manifold, there are three relevant operators and only one irrelevant operator, all lying in the four-dimensional space normal to the tangent space of two marginal operators. That is
$n_{IRR}=1,\,\text{and}\,\,n_{R}=3.$ (83)
@@
Overall three operators must be fine-tuned to the fixed point for the system to exhibit quantum critical behavior in the infrared limit. Hence, this conformal manifold is infrared-unstable. See appendix D for the exact form of the relevant and irrelevant operators and their respective scaling dimensions at a few selected points on the conformal manifold.
VIII.5 Phase boundary
@@
Recall that one of the conformal manifolds defined by $\hat{V}^{c}=\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$ is infrared stable. It has only one relevant operator, as shown in Eqn.82 (also in Table.4). Therefore, this conformal manifold is naturally embedded within the phase boundary of the gapless state. Eqn.82 suggests there are three irrelevant operators ($n_{IRR}=3$) at any point on the manifold. Following the spirit of the analysis in the $\mathcal{N}=2$ case, we can illustrate that locally the phase boundary shall be made up of this 3-dimensional subspace of irrelevant flows plus the 2-dimensional 2-sphere conformal manifold. To summarize, we shall have a five-dimensional space as the phase boundary. The sole relevant operator is directed orthogonal to this 5-dimensional phase boundary.
@@
VIII.6 Effective Yukawa field theory for the 5-dimensional phase boundary
In the previous section, where we studied the quantum criticality of the $\mathcal{N}=2$ phase, we developed an effective single boson Yukawa field theory for the conformal manifold of the shape of a ring in the three-dimensional parameter space (see section VIIC). We also found that the theory works well for the phase boundary of the gapless surface state. The two main ingredients that helped us develop this EFT are 1) the interaction matrix at the ring manifold is factorizable, and 2) there is only one relevant direction out of the conformal manifold that could be associated with the boson mass.
We observe that the phase boundary of the $\mathcal{N}=3$ flavor TI surface also satisfies the two ingredients. First, the 2-sphere conformal manifold embedded within the phase boundary is also factorizable. It is evident from its expression, $\hat{V}_{c}=\epsilon\,\tilde{e}^{3}_{\theta,\phi}\otimes\tilde{e}^{3}_{\theta,\phi}$, already in a tensor product form. Second, Eqn.82 suggests that the manifold has only one relevant operator. Therefore, the phase boundary of the $\mathcal{N}=3$ flavor TI surface near the infrared stable conformal manifold can also be expressed by a single boson Yukawa field theory in the infrared limit.
@@ VIII.6 Effective Yukawa field theory for the 5-dimensional phase boundary
Figure 9: RG flows of the parameters $V_{0}$, $V_{z}$ and $V_{x}$ when three-loop effects are taken into account for an $\mathcal{N}=2$ TI surface. The smooth conformal manifold is defined by the Eqn.89. The perfect ring manifold shown in Fig.7 is deformed but the closed loop structure remains intact. The definitions of other RG flow lines are the same as in fig.7.
IX Effects of $N_{L}$-loops, $N_{L}=2,3,4,…$
Further inclusion of effects of multiple-loop diagrams in the renormalization group equations can lead to the breaking of the emergent $SO(\mathcal{N})$ symmetry in the case of $\mathcal{N}=2,3$ studied in Section VII,VIII. The number of loops $N_{L}$ is related to the number of four fermion vertices $N_{v}$. $N_{v}=N_{L}+1$. In the context of the $\epsilon$-expansion, these contributions with multiple loops are also represented as higher-order effects so that symmetry is only weakly broken. For instance, the $N_{L}$-loop effects are proportional to $\epsilon^{L+1}$, $L\geq 1$. For two-loop effects, the contributions are of the order of $\epsilon^{3}$, while for three-loop effects, the contributions are of the order of $\epsilon^{4}$.
@@ IX Effects of $N_{L}$-loops, $N_{L}=2,3,4,…$
where $\beta(\hat{V})$ represents all the one-loop effects that already have been studied while $\beta_{M}(\hat{V})$ includes other multiple-loop effects that were not included in the studies in section VII.
@@
IX.1 A three-loop effect of Type A: Distorted Conformal Manifolds
@@
Let us see explicitly if the one-dimensional conformal manifold exists when the three-loop effects are included in the RG equation. Let us express the interaction matrix $\hat{V}$ in terms of Pauli matrices as shown in Eqn.32. The RG equation is then (See fig.6 for the Feynman diagram),
@@
Table 5: Table showing the list of fixed points and their stabilities for the $\mathcal{N}=2$ flavor surface when two-loop effects are included. Also, $\epsilon=D-2$ where D is the spacetime dimensions.
@@
IX.2 A two-loop effect of type B: Inducing flow in the Manifold
@@
Upon including the field renormalization effects, the RGE expressed in terms of $V_{0},V_{x},V_{z}$(see Eqn.32) takes the form,
@@
where $a=\frac{1}{32}$ and $\epsilon=D-2$.
IX.2.1 The fate of the conformal manifolds
@@
Taking into account the last two equations in Eqn.92, one can show that a finite flow is induced around the ring given by,
@@
Table.5 lists the fixed points of the RGE in Eqn.92 . Fig.10 shows the RG flow lines of the solutions to Eqn.92. Appendix.E lists the relevant and irrelevant operators and their scaling dimensions.